# Iterative Refinement of Project-Level Code Context for Precise Code Generation with Compiler Feedback

### v2

@@ 4 Experimental Setup
4.1 Datasets
To validate the effectiveness of ProCoder, we conducted experiments using the CoderEval benchmark Yu et al. (2023), a recently developed benchmark designed to evaluate models within realistic software development scenarios. Without loss of generalizability, we concentrate on the Python programming language within this dataset. This benchmark categories 230 test samples into six levels of context dependency: 1) self-contained: built-in types/functions, no imports required; 2) slib-runnable: standard libraries/modules, no installation needed; 3) plib-runnable: publicly-available libraries on PyPI/Maven; 4) class-runnable: code outside the function but within a class; 5) file-runnable: code outside the class but within the file; and 6) project-runnable: code in other source files. We concentrate on the last three dependency types, where the solutions are dependent on project-specific contexts. There are 55, 68, and 23 tasks associated with each dependency level, respectively.
@@ 4.2 Baseline Methods
ProCoder can function seamlessly be integrated into existing LLMs, requiring only black-box access to these models. In this paper, we select two state-of-the-art LLMs for code generation, namely GPT-3.5-Turbo OpenAI (2023a) and Code Llama (13B) Rozière et al. (2023), as our base models. To validate the effectiveness of ProCoder, we compare it with the following baselines:
@@
$\triangleright$ ReACC Lu et al. (2022). We employ the retrieval-augmented generation technique introduced in this baseline for code generation tasks. More precisely, we retrieve semantic contexts aligned with the task instructions and leverage them to augment the prompts of LLMs for better code generation.
@@
4.3 Implementation Details
We have selected GPT-3.5-turbo and Code-Llama 13B for investigation. For GPT-3.5-turbo, the invocation is performed through an online API. For Code-Llama 13B, we utilize its base variant, specifically designed for general code synthesis and understanding. [^2] The model can be accessed via https://huggingface.co/codellama/CodeLlama-13b-hf.
In the inference process, we set the decoding temperature to $0.7$, and adopt a top-$k$ sampling strategy. We implement the retrieval modules based on the text-ada model introduced by OpenAI OpenAI (2023b), which is effective in both natural language search and code search. The dimension for each embedding is 1,536. We retrieve at most 5 entries for each query. All the experiments in this paper are conducted on a Linux server with 128GB memory, with a single 32GB Tesla V100 GPU.
@@ 4.4 Evaluation Metrics
$\text{Pass}@k=\mathbb{E}\left[1-\frac{\binom{n-c}{k}}{\binom{n}{k}}\right]\,.$ (5)
@@ 5 Results and Analysis
5.1 Overall Performance of ProCoder
Table 2 shows the overall performance of ProCoder based on two LLMs, i.e., GPT-3.5-Turbo and Code Llama (13B), assessed against various baselines across different data splits. This figure shows that the ProCoder can significantly outperform other baselines on the project-level code generation across different data splits. This trend persists, with a few exceptions noted specifically in terms of Pass@$1$. We attribute such exceptions to variations in the generated solutions. Selecting a significantly larger $n$ (e.g., $1000$), as discussed in  Li et al. (2022), stabilizes expectations. Significantly, according to Code Llama (13B), when evaluated on the intricate project-runnable dataset, ProCoder demonstrates superior performance compared to the state-of-the-art RepoCoder, achieving an 87.7% higher success rate at Pass@$5$ and a 79.9% improvement at Pass@$10$. Moreover, it becomes evident that models incorporating contextual information, such as ReACC, RepoCoder, and ProCoder, exhibit a noteworthy performance superiority over the vanilla model, thereby affirming the practical value of contextual information.
Table 3: Pass Rates on CoderEval-Python of each approach with or without compiler feedback.
@@ 5.1 Overall Performance of ProCoder
Direct 20.65 26.66 29.13
Direct+CF 32.34 38.43 40.43
ReACC 34.13 41.44 43.48
ReACC+CF 36.60 44.05 46.52
RepoCoder 36.82 40.73 42.17
RepoCoder+CF 38.00 45.38 48.26
Figure 5: Performance of ProCoder, RepoCoder, and No Feedback baseline across three dependency levels.
5.2 Usefulness of Compiler Feedback
Here, we examine the usefulness of compiler feedback by integrating it into three baseline models: Vanilla, ReACC, and RepoCoder, each considered separately. We conducted experiments using the GPT-3.5-Turbo and subsequently reported the average score across the entire CoderEval dataset, as shown in Table 3. The table provides clear evidence that incorporating compiler feedback yields a notable enhancement in model performance for code generation. Specifically, a comparison between RepoCoder with and without compiler feedback reveals a substantial increase in Pass@$1$ from 36.82 to 38.00. More analytic experiments including the ablation study, ProCoder’s performance on function-level code generation, efficiency of static analysis and SQL queries, could be found in Appendix D
Figure 6: Compilation errors reported and fixed per iteration of ProCoder, RepoCoder, and No Feedback baselines.
5.3 Effectiveness of the Iterative Refinement
Here, we investigate the effectiveness of iterative refinement in code generation with compiler feedback. We conduct an ablation analysis on both RepoCoder and ProCoder, via removing or retaining the iterative refinement process. Figure 5 shows the performance of RepoCoder and ProCoder, with respect to varying iterations, on different data splits. This figure clearly illustrates that as the number of iterations increases, the performance of ProCoder also exhibits a corresponding improvement. It substantiates the efficacy of our suggested iterative refinement process, demonstrating its ability to enhance the generated code through multiple iterations progressively.
@@ 5.4 Error Analysis and Case Study
We also perform an error analysis of the generated code in iterative generation. We follow the categorization of errors defined in Section 2.2, examples of each error type can be found in Table 1. We exclude FUNC errors because they are reported in execution instead of by the compiler. Figure 6 shows the distribution of errors resolved iteratively by our ProCoder and two baselines. From this figure, we can see that the errors of various types can be effectively resolved after a single iteration of refinement. For instance, UNDEF errors were notably reduced from 5133 to 1042 after one iteration. Additionally, it was observed that the RepoCoder baseline, which operates without compiler feedback, managed to rectify API and syntax errors, corroborating the findings in  Zhang et al. (2023a). Nonetheless, it proved ineffective against UNDEF and OBJECT errors, likely due to the model’s lack of awareness regarding these errors in the absence of compiler feedback.
Figure 7: An error case from CoderEval-Python.
To thoroughly evaluate ProCoder’s effectiveness, we focused on scenarios where compilation was successful, but execution failed. In the case illustrated in Figure 7, ProCoder incorrectly excludes the microseconds field and erroneously adds a month field. This error stems from ambiguously stated task requirements and the model’s lack of familiarity with the Structure class’s format, despite its definition being available, resulting in misinterpretation of the intended functionality. Further examples are detailed in Appendix A. This observation inspires us to integrate a comprehensive reference comprising both documentation and web search results to provide explicit usage guidelines in our future work.
@@ 6 Related Work
Automated code generation has a history spanning several decades, with initial endeavors utilizing rule-based systems Backus et al. (1957); Woods (1973) and structured prediction Zelle and Mooney (1996); Zettlemoyer and Collins (2005). In recent years, the development of large-scale language models has led to the emergence of many prominent models in coding tasks. These include open-access models such as DeepSeek Coder Bi et al. (2024), Code Llama Rozière et al. (2023), and StarCoder Li et al. (2023), alongside commercial offerings like GPT-3.5 OpenAI (2023a), and GitHub Copilot Microsoft (2024) . These models and tools have demonstrated significant promise in enhancing code generation capabilities.
@@
Generating accurate code within a project poses challenges due to the modular design of software engineering, which results in cross-file dependency patterns Parnas (1972). Early work augmented N-gram, RNN, and LSTM models with an additional cache model to track local changes Tu et al. (2014); Hellendoorn and Devanbu (2017). Zhang et al. (2021) incorporate context from parent class to generate comments for the child class. Pashakhanloo et al. (2022) transforms projects into a relational database and proposed a graph walking method to traverse relation path and incorporate all relevant attributes along the way. Lyu et al. (2021) constructs an API dependency graph and utilizes a sequence-to-sequence model that incorporates graph embedding to predict the global-context-aware target code. Zan et al. (2022) first use private API documentation to improve code generation, and Zan et al. (2023) investigates how various components of API documentation influence the prediction accuracy. Shrivastava et al. (2023) proposes a code completion framework using repository-level prompt proposals, and then applying a classifier to classify useful proposals. Lu et al. (2022); Zhang et al. (2023a) propose to use single or multiple levels of retrieval-augmented generation mechanisms for code generation. Liao et al. (2023) proposes A3-CodGen, which utilize local, global, and third-party-library information for better context retrieval. Yu et al. (2023); Liu et al. (2023); Zhou et al. (2022); Ding et al. (2024) proposes benchmarks and datasets for repository-level code synthesis tasks.
Post-processing of LLMs for Code Generation
To identify correct code generated by LLM, researchers employ post-processing techniques to further rank and filter the generated code.  Inala et al. (2022) trained a fault-aware neural ranker that ranks multiple code samples based on compilation feedback. AlphaCode Li et al. (2022) and Shi et al. Shi et al. (2022) employ clustering and filtering methods based on the execution output of the generated program. Tsai et al. (2023) proposes to retrieve compilation error patches from a database of human expert guidance documents for RTL assembly.  Zhang et al. (2023c) rerank LLM outputs based on back translation. SelfEdit Zhang et al. (2023b) employs a generate-and-edit approach that utilizes execution results to improve the competitive programming task code quality.  Chen et al. (2023) propose a rubber duck debugging mechanism without human feedback, which identifies mistakes by investigating the execution results and explaining the generated code in natural language.
@@ 7 Conclusion
In this work, we show how to use project-specific context information, indexed structural and semantic, to fix compilation errors generated by compilers and improve the quality of code generated by LMs. Our experimental results show the increased prevalence of errors related to project contexts in project-level code generation compared to function-level code generation. The presented ProCoder can effectively fix the compilation errors by retrieving related context from the project, thus significantly improving the native LLM baselines on over 80% relative pass rates in generating functions dependent on project-specific contexts.
@@
In this study, we utilize compilation information as a means to validate programs. However, it is important to note that even programs that compile successfully can experience execution failures. Furthermore, the successful compilation of programs does not guarantee their safety for execution. As a result, while our findings indicate improvements in quality metrics through the correction of code to properly leverage context, it is imperative to undertake additional verification methods such as testing and manual review to ascertain the functional correctness of the generated code. The challenge of ensuring functional correctness of code encompasses various aspects, including compliance with task requirements, adherence to pre/post-conditions and security requirements, and preserving robustness in generating code. With many questions unanswered, we hope our study can promote a broader view of utilizing computational linguistic technologies in the realm of automated software engineering.
@@
We meticulously ensured that all code and models integrated into our research adhere to open access policies as outlined by the Creative Commons license. The methodology ensures full compliance with copyright and intellectual property laws, thereby eliminating any potential for infringement or unauthorized use of protected materials. By exclusively utilizing resources that are freely available and legally distributable, we maintain the highest standards of ethical conduct in research. This approach fosters an environment of transparency and respect for the intellectual property rights of others. Our commitment to these principles ensures that our work advances the frontiers of knowledge in a manner that is both legally sound and ethically responsible.
Appendix A More details about the Error Cases
In this section, we delve into additional cases where ProCoder struggles to generate viable solutions, elucidating the underlying causes to foster further research. We identify two primary factors leading to execution errors despite successful compilation: degenerate solutions, such as overly simplistic or redundant code. This phenomenon is discovered and explored in Zhang et al. (2023c). Figure A1 exemplifies this issue, showcasing a solution that omits necessary validity checks and error handling, neither of which are specified in the task requirements, nor produced by ProCoder’s outputs. Our analysis reveals that degenerate solutions frequently pass compilation, rendering compiler-based verification ineffective.
Another prevalent mistake involves misinterpreting task requirements, resulting in solutions that lack logical coherence. Figure A2 depicts an instance where the assignment was to leverage the pre-existing Point class; instead, the language model disregards this specification and redundantly recreates the class. This underscores the challenge large language models (LLMs) face in accurately comprehending prompts and generating appropriate solutions.
Figure A1: An error case of degenrate solution.
Figure A2: An error case of misinterpreting task requirements.

### v3

@@ 4 Experimental Setup
4.1 Models and Datasets
To validate the effectiveness of CoCoGen, we select GPT-3.5-Turbo and Code-Llama 13B basehttps://huggingface.co/codellama/CodeLlama-13b-hf language models for investigation. The technical details of models and their invocation for inference are presented in Appendix A.
We conduct experiments using the Python split of the CoderEval benchmark Yu et al. (2024), referred to as CoderEval-Python. It is a benchmark designed to evaluate models within realistic software development scenarios. Without loss of generalizability, we concentrate on the Python programming language within this dataset. This benchmark categorizes 230 test samples into six levels of context dependency: 1) self-contained: built-in types/functions, no imports required; 2) slib-runnable: standard libraries/modules, no installation needed; 3) plib-runnable: publicly-available libraries on PyPI/Maven; 4) class-runnable: code outside the function but within a class; 5) file-runnable: code outside the class but within the file; and 6) project-runnable: code in other source files. We concentrate on the last three dependency types, where the solutions are dependent on project-specific contexts. There are 55, 68, and 23 tasks associated with each dependency level, respectively.
We also evaluate CoCoGen on two function-level code generation benchmarks, namely HumanEval Chen et al. (2021) and MBPP Austin et al. (2021), as well as a project-level code completion benchmark named CrossCodeEval Ding et al. (2023), to further validate its generalizability across other coding tasks. the statistics and experimental results of these benchmarks can be found in Appendix B and Appendix F.
@@ 4.2 Baseline Methods
CoCoGen can function seamlessly and be integrated into existing LLMs, requiring only black-box access to these models. In this paper, we select two state-of-the-art LLMs for code generation, namely GPT-3.5-Turbo OpenAI (2023) and Code Llama (13B) Rozière et al. (2023), as our base models. To validate the effectiveness of CoCoGen, we compare it with the following baselines:
@@
$\triangleright$ ReACC Lu et al. (2022). We employ the retrieval-augmented generation technique introduced in this baseline for code generation tasks. More precisely, we retrieve project contexts aligned with the task instructions semantically through embedding similarity, and leverage them to augment the prompts of LLMs for better code generation.
@@ 4.3 Evaluation Metrics
We employ the Pass@$k$ metric Chen et al. (2021); Yu et al. (2024) for code generation tasks, and employ the code exact match (C-EM), code edit similarity (C-ES), identifier exact match (I-EM), and identifier F1 score (I-F1) to evaluate both the code match rate and the identifier match rate for code completion tasks, follows Ding et al. (2023). We present the details of the Pass@$k$ metric for code generation. For metrics used in the code completion task, readers are referred to Appendix C.
Pass@$k$.
@@
$\text{Pass}@k=\mathbb{E}\left[1-\frac{\binom{n-c}{k}}{\binom{n}{k}}\right]$ (5)
@@ 5 Results and Analysis
Figure 5: Pass@10 of CoCoGen, RepoCoder, and No Feedback baseline across three dependency levels
5.1 Overall Performance of CoCoGen
Table 2 shows the overall performance of CoCoGen, assessed against various baselines, on the CoderEval Yu et al. (2024) and the CrossCodeEval Ding et al. (2023) dataset, respectively. This table shows that the CoCoGen can significantly outperform other baselines on the project-level code generation task. This trend persists, with a few exceptions noted specifically in terms of Pass@$1$. We attribute such exceptions to variations in the generated solutions. Selecting a significantly larger $n$ (e.g., $1000$), as discussed in  Li et al. (2022), stabilizes the result and eliminates these expectations. Moreover, it becomes evident that models incorporating contextual information, such as ReACC, RepoCoder, and CoCoGen, exhibit a noteworthy performance superiority over the no context (i.e., direct) model, thereby affirming the practical value of the context in project-level code generation.
5.2 Effectiveness of the Iterative Refinement
Here, we investigate the effectiveness of iterative refinement in code generation with compiler feedback. We conduct an ablation analysis on both RepoCoder and CoCoGen, via removing or retaining the iterative refinement process. Figure 5 shows the performance of RepoCoder and CoCoGen, with respect to varying iterations, on different data splits. This figure clearly illustrates that as the number of iterations increases, the performance of CoCoGen also exhibits a corresponding improvement. The improvement substantiates the efficacy of our suggested iterative refinement process, demonstrating its ability to enhance the generated code through multiple iterations progressively.
Table 3: Pass rates of CoCoGen with components ablated, based on GPT-3.5-Turbo model using the CoderEval-Python dataset
@@ 5.2 Effectiveness of the Iterative Refinement
CoCoGen 28.01 43.01 46.58
- w/o CF and SQL (RepoCoder) 28.72 34.44 36.30
- w/ CF, w/o SQL, w/o Semantic 25.69 37.50 41.78
- w/ CF and SQL, w/o Semantic 26.37 38.31 41.78
- w/ CF and Semantic, w/o SQL 27.39 40.02 44.45
5.3 Ablation on Components
To explore the impact of each component within CoCoGen on overall performance, we remove various components from CoCoGen, including the compiler, SQL retriever, and semantic retriever. Notably, the configuration that excludes the compiler and SQL, relying solely on semantic retrieval, is known as RepoCoder Zhang et al. (2023a).
Table 3 demonstrates the pass rates of CoCoGen across CoderEval-Python. The data indicates that incorporating compiler feedback does improve accuracy, although not markedly substantial. The reason is that highlighting compilation errors alone does not provide related context for resolving them, which is important in project-level coding problems. Also, compared to RepoCoder—which relies solely on semantic retrieval, integrating compiler feedback with project-specific context results in a stable performance improvement. Results on each data split can be found in Appendix F.2.
Further analytical experiments, including the evaluation of benefits from solely compiler feedback, the performance of CoCoGen on project-level code completion and function-level code generation, and the efficiency of static analysis and SQL queries, can be found in Appendix F.
Figure 6: Compilation errors fixed per iteration of CoCoGen, RepoCoder, and No Feedback baselines
@@ 5.4 Error Analysis and Case Study
We also perform an error analysis of the generated code in iterative generation. We follow the categorization of errors defined in Section 2.2, examples of each error type can be found in Table 1. Figure 6 shows the distribution of errors resolved iteratively by our CoCoGen and two baselines. From this figure, we can see that the errors of various types can be effectively resolved after a single iteration of refinement. For instance, UNDEF errors are notably reduced from 5,133 to 1,042 after one iteration. Additionally, it is observed that the RepoCoder baseline, which operates without compiler feedback, manages to rectify API and syntax errors, corroborating the findings in  Zhang et al. (2023a). Nonetheless, RepoCoder proves ineffective against UNDEF and OBJECT errors, likely due to the model’s lack of awareness regarding these errors in the absence of compiler feedback.
Figure 7: An example of a runtime error in CoCoGen’s generated code for a CoderEval test case
To thoroughly evaluate CoCoGen’s effectiveness, we focus on scenarios where compilation is successful, but execution fails. In the case illustrated in Figure 7, CoCoGen incorrectly excludes the microseconds field and erroneously adds a month field. This error stems from ambiguously stated task requirements and the model’s lack of familiarity with the Structure class’s format, despite its definition being available, resulting in misinterpretation of the intended functionality. Further examples are detailed in Appendix E. The observations inspire us to integrate a comprehensive reference comprising documentation, web search results, and code execution log to provide explicit guidelines for code generation in our future work.
@@ 6 Related Work
Automated code generation has a history spanning several decades, with initial endeavors utilizing rule-based systems Woods (1973) and structured prediction Zelle and Mooney (1996); Zettlemoyer and Collins (2005). In recent years, the development of LLMs has led to the emergence of many prominent models in coding tasks. These include open-access models such as DeepSeek Coder Bi et al. (2024), Code Llama Rozière et al. (2023), and StarCoder Li et al. (2023), alongside commercial offerings like GPT-3.5 OpenAI (2023) and GitHub Copilot Microsoft (2024). These models and tools have demonstrated significant promise in enhancing code generation capabilities.
@@
Generating accurate code within a project poses challenges due to the modular design of software engineering, which results in cross-file dependency patterns Parnas (1972). Early works augmented N-gram, RNN, and LSTM models with an additional cache model to track project-level changes Tu et al. (2014); Hellendoorn and Devanbu (2017). Pashakhanloo et al. (2023) transformed projects into a relational database and proposes a graph walking method to traverse this database. Zan et al. (2022) first used private API documentation to improve code generation, and Zan et al. (2023) investigated how various components of API documentation influence prediction accuracy. Shrivastava et al. (2023) proposed a code completion framework using a classifier to filter useful repository-level prompt proposals. Lu et al. (2022); Zhang et al. (2023a) proposed to use single or multiple levels of retrieval-augmented generation mechanisms for code generation. Liao et al. (2023) proposed A3-CodGen, which utilizes local, global, and third-party-library information for better context retrieval. Yu et al. (2024); Liu et al. (2023); Zhou et al. (2022); Ding et al. (2023) proposed benchmarks and datasets for repository-level code generation tasks.
Post-processing of LLMs for Code Generation.
To identify correct code generated by LLM, researchers employ post-processing techniques to further rank and filter the generated code. Inala et al. (2022) trained a fault-aware neural ranker that ranks multiple code samples based on compilation feedback. AlphaCode Li et al. (2022) and Shi et al. (2022) employed filtering methods based on the execution feedback. Zhang et al. (2023c) reranked LLM outputs based on back translation. SelfEdit Zhang et al. (2023b) employed a generate-and-edit approach that utilizes execution results to improve the competitive programming task code quality. Chen et al. (2023) proposed a rubber duck debugging mechanism without human feedback, which identifies mistakes in generated code by investigating the execution results, and explaining the generated code in natural language.
@@ 7 Conclusion
In this paper, we show how to use project-specific context information, indexed structural and semantic, to fix compilation errors generated by compilers and improve the quality of code generated by LLMs. Our experimental results show the increased prevalence of errors related to project contexts in project-level code generation compared to function-level code generation. The presented CoCoGen can effectively fix the compilation errors by retrieving related context from the project, thus significantly improving the native LLM baselines on over 80% relative pass rates in generating functions dependent on project-specific contexts.
Acknowledgements
This work is supported by the Major Program (JD) of Hubei Province (Grant No. 2023BAA024), and the National Natural Science Foundation of China under grant No. 62102157. This work is also partially supported by Huawei. Fangxin Lu is a visiting student at Huazhong University of Science and Technology, and she is from South-Central Minzu University for Nationalities.
@@ Acknowledgements
In this paper, we utilize compilation information as a means to validate programs. However, it is important to note that even programs that compile successfully can experience execution failures. Furthermore, the successful compilation of programs does not guarantee their safety for execution. As a result, while our findings indicate improvements in quality metrics through the correction of code to properly leverage context, it is imperative to undertake additional verification methods such as testing and manual review to ascertain the functional correctness of the generated code. The challenge of ensuring functional correctness of code encompasses various aspects, including compliance with task requirements, adherence to pre/post-conditions and security requirements, and preserving robustness in generating code. With many questions unanswered, we hope our study can promote a broader view of utilizing computational linguistic technologies in the realm of automated software engineering.
@@
We meticulously ensure that all code and models integrated into our research adhere to open-access policies as outlined by the Creative Commons license. The methodology ensures full compliance with copyright and intellectual property laws, thereby eliminating any potential for infringement or unauthorized use of protected materials. By exclusively utilizing resources that are freely available and legally distributable, we maintain the highest standards of ethical conduct in research. This approach fosters an environment of transparency and respect for the intellectual property rights of others. Our commitment to these principles ensures that our work advances the frontiers of knowledge in a manner that is both legally sound and ethically responsible.
Appendix A More Details on Investigated Large Language Models
In this paper, we have selected GPT-3.5-Turbo and Code-Llama 13B for investigation.
GPT-3.5-Turbo OpenAI (2023).
It is a large-scale decoder-only model based on the Transformer architecture Vaswani et al. (2017). It is pre-trained on a diverse array of data, encompassing both natural language and code, and can learn a specific task given an instruction and several demonstration examples. We perform the inference by invocating the model through an online API.
Code Llama Rozière et al. (2023).
It is a family of LLMs for code based on Llama 2 Touvron et al. (2023) and further trained on 1TB code and natural language tokens. It can be used for code completion and generation following natural language instructions. For Code-Llama 13B, we utilize its base variant, specifically designed for general code generation and understanding.The model can be accessed via https://huggingface.co/codellama/CodeLlama-13b-hf.
In the inference stage, we set the decoding temperature to $0.7$, and adopt the top-$k$ sampling strategy. We implement the retrieval modules based on the text-ada model introduced by OpenAI OpenAI (2022), which is effective in both natural language search and code search. The embedding dimension is 1,536 for the text-ada model. We retrieve at most 5 entries for each query. All the experiments in this paper are conducted on a Linux server with 128GB memory, with four 32GB Tesla V100 GPUs.
Appendix B More Details on the Investigated Dataset
In this paper, we also use three recently proposed benchmarks evaluating the project-level code completion and function-level code generation ability. We provide the details of each benchmark.
CrossCodeEval Ding et al. (2023).
It is a project-level code completion benchmark that necessitates cross-file contextual understanding to complete the code accurately. CrossCodeEval is built on a diverse set of real-world, open-sourced, permissively-licensed repositories, including 471 repositories, 2,665 test cases across 1,368 files for Python language Ding et al. (2023).
HumanEval Chen et al. (2021).
It is a function-level code generation benchmark that consists of 164 programming problems with corresponding Python solutions. These problems cover a range of difficulty levels and programming concepts. HumanEval is widely used in evaluating LMs for code generation tasks due to its focus on real-world coding scenarios and its comprehensive test cases.
MBPP Austin et al. (2021).
It is a function-level code generation benchmark consisting of 974 crowd-sourced Python programming problems. These problems are designed to be solvable by entry-level programmers and cover programming fundamentals and standard library functionalities. Each problem includes a task description, a code solution, and several automated test cases.
Appendix C More Details on the Evaluation Metrics
In this work, we use several evaluation metrics for method generation and line completion tasks. This includes:
Exact Match.
This metric assesses whether each character of the model’s predicted code exactly matches each character of the correct answer. All characters are the same yields $EM=1$, while any discrepancy results in $EM=0$. This is a strict all-or-nothing metric, where a single character difference results in a score of 0. The average value of EM on all test cases is reported.
Edit Similarity Levenshtein (1965).
This metric quantifies how dissimilar two strings are to one another. It is measured by counting the minimum number of operations required to transform one string into the other:
$\operatorname{lev}(a,b)=\begin{cases}|a|&\text{ if }|b|=0,\\ |b|&\text{ if }|a|=0,\\ \operatorname{lev}(\operatorname{tail}(a),\operatorname{tail}(b))&\text{ if }\operatorname{head}(a)=\operatorname{head}(b),\\ 1+\min\begin{cases}\operatorname{lev}(\operatorname{tail}(a),b)\\ \operatorname{lev}(a,\operatorname{tail}(b))\\ \operatorname{lev}(\operatorname{tail}(a),\operatorname{tail}(b))\end{cases}&\text{ otherwise }\end{cases}$
F1 Score Schütze et al. (2008).
This metric evaluates the balance between precision and recall for the retrieved identifiers. The F1 score is based on the number of identifiers that both the prediction and the ground truth share. Precision is defined as the ratio of the shared identifiers to the total number of identifiers in the prediction, while recall is the ratio of the shared identifiers to the total number of identifiers in the ground truth. The F1 score is then calculated as the harmonic mean of the recall rate and the precision rate:
$F_{1}=\frac{2}{\mathrm{recall}^{-1}+\mathrm{precision}^{-1}}$ (7)